The Springfield Company has three service departments, S1, S2, and S3, and two production departments, P1 and P2. The following data relate to Springfield’s allocation of service department costs:
Budgeted Costs Number of Employees
S1 $4,588,000 92
S2 2,834,000 64
S3 2,760,000 41
P1 196
P2 294

Service department costs are allocated by the direct method. The number of employees is used as the allocation base for all service department costs.

Collapse question part

(a)

Allocate service department costs to production departments.
Service Service dept. Cost Allocated to
Department Costs P1 P2
S1 $4,588,000 $

$

S2 2,834,000 $

S3 2,760,000 $
Answer #1

a.

Number of employees in P1 = 196

Number of employees in P2 = 294

Total number of employees in P1 and P2 = Number of employees in P1 + Number of employees in P2

= 196 + 294

= 490

Service department Service department cost Cost allocated to
P1 P2
S1 4,588,000 4,588,000 x 196/490 = $1,835,200 4,588,000 x 294/490 = $2,752,800
S2
2,834,000
2,834,000 x 196/490 = $1,133,600 2,834,000 x 294/490 = $1,700,400
S3 2,760,000 2,760,000 x 196/490 = $1,104,000 2,760,000 x 294/490 = $1,656,000


kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.